OpenStudy (.whitedragon.):

I got 152 What is the third quartile (Q3) of the data set? 165, 85, 105, 105, 152, 144, 140, 160, 80 A. 160 B. 156 C. 105 D. 140

OpenStudy (anonymous):

First, list in order 80, 85, 105, 105, 140, 144, 152, 160, 165 Q1 = {80.85} 2/9 = .22222 < .25 Q2 = {105,105) 4/9 = .4444 < .5 Q3 = {140,144} 6/9 = .6666 < .75 Q4 = {152, 160, 165} 9/9 = 1 = 1

OpenStudy (anonymous):

So you were close.... what do you think the answer would be? :))

OpenStudy (.whitedragon.):

maybe 140 then ?

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

I have to go class is starting!

OpenStudy (anonymous):

its 156 guys, sorry for butting in

OpenStudy (anonymous):

u look for it after the median, so u have 4 numbers to consider 144, 152, 160, 165 the 3rd quartile =( 152+160) /2 = 156

OpenStudy (anonymous):

@.WhiteDragon.